OpenStudy (anonymous):

find the absolute extrema of the function y=x^3-(3/2)x^2, [-1,2]

OpenStudy (anonymous):

My answeres are f(-1)= -5/2 f(0)=0 f(1)= -1/2 f(2)=2 I just dont know which is the min and max

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

ok. got it

OpenStudy (anonymous):

so are the mins and max the endpoints or the hills and valleys??